find angle A if angle A is complementary to angle B, angle B is supplementary to angle C, angle B= 15x-2 and angle C= 25+22
Answer by cleomenius(959)

I am assuming you stated the problem incorrectly and angle B should be 25x + 22.
15x -2 + 25x + 22 = 180
40x = 160
x = 4
15(4) -2 = 58 degrees angle B
25(4) + 22 = 122 degrees angle C
122 + 58 = 180 degrees.
Angle A, the complement of angle B is 32 degrees.
Cleomenius.